OpenStudy (anonymous):

What is (1.89e-2)/(2.7e-5) ??? I can break down the more simple ones but this one is killing me

OpenStudy (anonymous):

I know the answer is supposed to be 700 aka 7e2 but i keep getting 5.103e3

OpenStudy (anonymous):

\[\frac{e^{-2}}{e^{-5}}=e^{-2+5}=e^3\]

OpenStudy (anonymous):

and \[\frac{1.89}{2.7}=.7\] so you get \[.7e^3\]

OpenStudy (anonymous):

unless of course that "e" is the "e" at the end of your calculator, in which case it is actually 10. is that what you mean?

OpenStudy (anonymous):

any easy way to divide that? I have an exam in 2 hours and we dont get a calc.

OpenStudy (anonymous):

if it is really a ten then you get \[.7\times 10^2=70\] but if it is the number e, then it is e. done

OpenStudy (anonymous):

no there is no easy way. you have to write \[\frac{1.89}{2.7}=\frac{18.9}{27}\] and then use long division

OpenStudy (anonymous):

ok...sounds good...perfect explaination...thank you!

OpenStudy (anonymous):

yw